40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Oct 30 19:20:27 2019, gautam, Update, SUS, PRM suspension issues PRMUL.pdfPRMnormal.pdfPRM-Sensors_noise.pdfPRMsuspensionWonky.png
    Reply  Wed Oct 30 23:12:27 2019, Koji, Update, SUS, PRM suspension issues 
Message ID: 15002     Entry time: Wed Oct 30 19:20:27 2019     Reply to this: 15003
Author: gautam 
Type: Update 
Category: SUS 
Subject: PRM suspension issues 

While I was trying to lock the PRMI this evening, I noticed that I couldn't move the REFL beamspot on the CCD field of view by adjusting the slow bias voltages to the PRM. Other suspensions controlled by c1susaux seem to respond okay so at first glance it isn't a problem with the Acromag. Looking at the OSEM sensor input levels, I noticed that UL is much lower than the others - see Attachment #1, seems to have happened ~100 days ago. I plugged the tester box in to check if the problem is with the electronics or if this is an actual shorting of some pins on the physical OSEM as we had in the past. So PRM watchdog is shutdown for now and there is no control of the optic available as the cables are detached. I will replace the connections later in the evening.

Update 10pm:

  1. Measured coil inductances with breakout board and LCR meter - all 5 coils returned ~3.28-3.32 mH.
  2. Measured coil resistances with breakout board and DMM - all 5 coils returned ~16-17 ohms.
  3. Checked OSEM PD capacitance (with no bias voltage) using the LCR meter - each PD returned ~1nF.
  4. Checked resistance between LED Cathode and Anode for all 5 LEDs using DMM - each returned Hi-Z.
  5. Checked resistance between PD Cathode and Anode for all 5 PDs using DMM - each returned ~430 kohms.
  6. Checked that I could change the slow bias voltages and see a response at the expected pins (with the suspension disconnected).

Since I couldn't find anything wrong, I plugged the suspension back in - and voila, the suspect UL PD voltage level came back to a level consistent with the others! See Attachment #2.

Anyway, I had some hours of data with the tester box plugged in - see Attachment #3 for a comparison of the shadow sensor readout with the tester box (all black traces) vs with the suspension plugged in, local damping loops active (coloured traces). The sensing noise re-injection will depend on the specifics of the  local damping loop shapes but I suspect it will limit feedforward subtraction possibilities at low frequencies.

However, I continue to have problems aligning the optic using the slow bias sliders (but the fast ones work just fine) - problem seems to be EPICS related. In Attachment #4, I show that even though I change the soft PITCH bias voltage adjust channel for the PRM, the linked channels which control the actual voltages to the coils take several seconds to show any response, and do so asynchronously. I tried restarting the modbus process on c1susaux, but the problem persists. Perhaps it needs a reboot of the computer and/or the acromag chassis? I note that the same problem exists for the BS and PRM suspensions, but not for ITMX or ITMY (didn't check the IMC optics). Perhaps a particular Acromag DAC unit is faulty / has issues with the internal subnet?

Attachment 1: PRMUL.pdf  391 kB  Uploaded Wed Oct 30 20:24:15 2019  | Hide | Hide all
PRMUL.pdf
Attachment 2: PRMnormal.pdf  608 kB  Uploaded Wed Oct 30 23:23:54 2019  | Hide | Hide all
PRMnormal.pdf
Attachment 3: PRM-Sensors_noise.pdf  102 kB  Uploaded Wed Oct 30 23:27:44 2019  | Hide | Hide all
PRM-Sensors_noise.pdf PRM-Sensors_noise.pdf
Attachment 4: PRMsuspensionWonky.png  17 kB  Uploaded Thu Oct 31 00:02:53 2019  | Hide | Hide all
PRMsuspensionWonky.png
ELOG V3.1.3-